800 volunteers produce a world-class operating system and give it away!

It may sound incredible, but it’s a fact. Not only that, but the fruit of their labours, Debian GNU/Linux, is more reliable, stable, secure and maintainable than practically anything you could buy. Debian’s package manager, APT, is often said to be the best package manager around — update your entire system with just two commands!

Over the past 8 years Debian has grown into the operating system of choice for many organisations and individuals alike. Debian has been run on everything from embedded devices to the most powerful super computers. Some of Debian’s larger users include NASA and France Telecom. Furthermore, Debian supports more processor architectures than any other Linux-based operating system around, and is working on support for even more.

Unlike most publicity of this sort, where you know there’s got to be a catch, with Debian there’s no reason for us to bend the truth. Since we don’t make any money from our users, there’s no real incentive to overstate the qualities of the product.

All Debian developers have undergone a stringent selection process to ensure that only the best people develop the best operating system. Most Debian developers work in their free time for no monetary reward, so you can trust that they are highly motivated, ensuring that their handiwork is nothing but perfect.

For this reason you can be assured that Debian is going to be actively developed for the foreseeable future. If you ever need technical support, then help is readily available right from the people who make it.

This is why, when you choose Debian, you know you’re in safe hands. And, don’t forget, Debian is absolutely free so you have nothing to lose.

If this has sparked your interest, then come to the .ORG stand (stand 230) and ask us what on earth we think we’re doing. Look out for people wearing Debian T-shirts (with our logo, see above), or visit the web site:

http://www.uk.debian.org/

If you’re not interested in this, please give it to someone who might be, since being a non-profit organisation means that we have to try to make the best use of the limited

funds available. Thanks.